Emily buys a pen and a pad of paper for �2.20. The paper costs �2 more than the pen. How much is the pen?
Answer by CubeyThePenguin(3113)

x = cost of pen
y = cost of pad of paper

x + y = 2.2
y = x + 2

x + (x + 2) = 2.2
2x + 2 = 2.2
x = 0.1

The pen costs $0.10.